A quantity of bolts are to be produced from a bar of 3/4" square stock, 8 1/4' in length. Allowing 1/8" for waste in cutting off each bolt, calculate number of bolts that can be made.
Description of bolt:
3/4" square head.
3/8" diameter of bolt cylinder.
2 5/8" total length of bolt.

My attempt at solution:
8 1/4' = 99"
2 5/8" - 1/8" = 2 4/8" = 2.5"
99 / 2.5 = 39.6 = 40 bolts.
Where am I correct ?

Answer by solver91311(24713)



You missed it by THAT much. If you are going to allow 1/8" waste for each bolt, then you must ADD 1/8" to the 2 5/8" required for the bolt to get the amount of metal consumed each time you cut off enough for one bolt.

So

8 1/4' = 99"
2 5/8" + 1/8" = 2 3/4" = 2.75"
99 / 2.75 = 36


John

My calculator said it, I believe it, that settles it